# When Does Meaning Backfire? Investigating the Role of AMRs in NLI

### v1

@@ Abstract
Natural Language Inference (NLI) relies heavily on adequately parsing the semantic content of the premise and hypothesis. In this work, we investigate whether adding semantic information in the form of an Abstract Meaning Representation (AMR) helps pretrained language models better generalize in NLI. Our experiments integrating AMR into NLI in both fine-tuning and prompting settings show that the presence of AMR in fine-tuning hinders model generalization while prompting with AMR leads to slight gains in GPT-4o. However, an ablation study reveals that the improvement comes from amplifying surface-level differences rather than aiding semantic reasoning. This amplification can mislead models to predict non-entailment even when the core meaning is preserved.
@@
[…]
@@ 1 Introduction
Natural language inference (NLI; Dagan et al., 2010) is a popular task in NLP where the solver is given a premise and a hypothesis, and asked to determine whether the hypothesis is true if the premise is true. The label space consists of three labels: entailment if the hypothesis is true, contradiction if the hypothesis is false, and neutral if the truth value of the hypothesis cannot be determined; this can also be condensed in two labels: entailment and non-entailment. As shown in Figure 1 “Athletes introduced the secretaries” should be entailed by “Serious athletes introduced the secretaries.” Therefore, the label should be entailment because the truth of the premise indicates truth of (or entails) the hypothesis.
As a meaning-focused task, NLI aligns well with the motivation behind AMRs, i.e., to abstract sentence meaning beyond surface form, given NLI models’ tendencies to adopt shallow heuristics rather than understanding the relationship between the premise and the hypothesis, leading to poor generalization to novel data (Gururangan et al., 2018; Poliak et al., 2018; McCoy et al., 2019; Serrano et al., 2023). In this paper, we investigate whether incorporating AMRs as additional input - either during (a) fine-tuning or (b) prompting - can encourage models to attend more to abstract meaning, thereby improving generalization and overall performance. As illustrated in Figure 1, we add AMRs to either the training data or prompts then evaluate how the addition of AMR affects generalization performance. We find that AMRs generally hinder performance in both fine-tuning and prompting settings, with the exception of prompting on HANS. However, this improvement appears to stem from AMRs amplifying surface-level differences rather than capturing deeper semantic meaning.
@@ 3.1 Data & Models
In these experiments, we use two datasets: MNLI (Williams et al., 2018) and HANS (McCoy et al., 2019). MNLI is a crowdsourced dataset, with a test set that is not available to the public. We follow prior work (Wang et al., 2018; Devlin et al., 2019) in taking one of its two developmental splits as the evaluation dataset. Specifically, we take the matched developmental set to use as our evaluation dataset. The training dataset includes 297k sentence pairs, while the evaluation set contains around 10k pairs. HANS is a template-based evaluation dataset, with 30k examples. Unlike MNLI and other NLI datasets, its label space consists of only two labels–entailment and non-entailment. We follow prior work (McCoy et al., 2020; Min et al., 2020) in collapsing the model’s neutral and contradiction predictions to the single non-entailment label when calculating evaluation metrics.
@@ 3.2 Experiment 1: Can fine-tuned models benefit from AMR in NLI?
We adopt the setup and hyperparameters of previous work in MNLI fine-tuning and HANS evaluation (McCoy et al., 2020; Min et al., 2020), detailed in Appendix A. We integrate AMR into the models as text, via linearization, removing all newlines and whitespace sequences longer than length two.
@@ 3.3 Experiment 2: Can prompt-based models benefit from AMR in NLI?
In this experiment, we evaluate whether incorporating AMRs improves LLMs’ performance on NLI, on both the MNLI and HANS dataset, the latter of which remains challenging even after fine-tuning. Jin et al. (2024) find that only instruction-tuned GPT models are capable of reliably processing AMRs. We therefore restrict our evaluation to GPT-4o in zero-shot and 5-shot settings. The template can be found in Appendix B.
@@ 4.1 Experiment 1
We report the accuracies of our fine-tuning models with and without AMRs in Table 1. We report numbers from prior work (McCoy et al., 2020; Min et al., 2020) in addition to our experiments to serve as comparison baselines and to ensure our setup is correct. Our reported numbers are an average across 10 runs with varying seed.
As shown in Table 1, AMR augmentation does not yield improvements in MNLI performance, nor HANS generalization. Perhaps analogously to previous data-driven attempts at improving generalization (Clark et al., 2019; Min et al., 2020; Yaghoobzadeh et al., 2021), additional AMR information as either text or graph does not affect MNLI performance. An analysis in to their confusion matrices reveals AMR adds or subtracts little in terms of MNLI label decision boundary. On HANS performance, We discuss two main findings.
@@
AMR-only models predict the entailment label for 98.3% of HANS examples, compared to the baseline models at 94.7%. We attribute this to an intensified version of the baseline models’ heuristic correlating overlap between the hypothesis and premise to the entailment label, dubbed the lexical overlap heuristic (McCoy et al., 2019). We note this is concurrent with a still competitive MNLI performance, at 84%. We discuss this phenomenon in more detail in Sections C.1 and C.2.
@@
On the other hand, combining AMR information with text strongly affects HANS label decision boundaries in the opposite direction, overriding various shallow heuristics that favor the entailment label observed in McCoy et al. (2020) and in our baseline and AMR-only experiments. Our +AMR as text models 86.6% of HANS examples, and +AMR as graph models 86.9%, even predicting non-entailment on highly overlapping examples. We attempt to disentangle the effects of AMRs and text in a combined representation in Section C.2, where we find that while AMR can be used to perform NLI, it is less effective than text input and combining the two introduces new artifacts that are more difficult to interpret.
@@ 4.2 Experiment 2
The results for prompting with GPT-4o are shown in Table 2. We report only the zero-shot results in the main text, as they yield similar overall performance and prediction patterns. Results for the five-shot setting are provided in Appendix F. Two main observations emerge.
@@
As shown in the table, model performance is consistently lowest when prompted with AMRs alone, while including the original sentence improves results. We find this is because AMRs lead models to make more negative predictions (see Appendix D). To test this statistically, we fit a logistic regression model predicting non-entailment using SMATCH++ (Opitz, 2023) between hypothesis and premise AMRs and data source (gold vs. predicted). A significant negative interaction ($\beta$ = -0.042, $p$ < 2e-16) shows that SMATCH similarity influences model predictions more than gold labels.
@@
Further analysis reveals that AMR’s sensitivity to surface-level lexical and syntactic variation leads to low structural overlap between semantically equivalent expressions, [^2] See Appendix E for an example. misleading the model toward non-entailment. This also explains why, on the HANS test set, prompts that include both the sentence and its AMR lead to the highest rate of negative predictions: the AMR representation amplifies subtle differences between two otherwise similar strings, making semantic mismatches more salient and pushing the model toward rejecting entailment. Such nuanced contrasts are what HANS is designed to probe in language models, prompting GPT-4o to overpredict non-entailment.
@@
As shown in Figure 2, when GPT-4o is prompted with both sentence and AMR inputs, accuracy slightly increases for inputs longer than 80 words. However, this performance remains lower than that of sentence-only prompts across most length bins. We find no evidence that AMR-only prompts enhance robustness to longer sequences.
@@ 4.3 Summary
Our LLM experiments showcases similar preference towards the non-entailment label. This suggests that AMRs effectively highlight subtle distinctions between minimal pairs, explaining improved HANS performance. However, for simpler examples, this heightened contrast can cause the model to overpredict No, even for entailing sentence pairs.
@@ 5 Conclusion
Future work could explore more diverse linearization formats, prompt designs, and integration strategies that align AMR structure more directly with model attention or reasoning processes. In addition to Python-like and natural language-like representation of AMR’s structure (Srivastava et al., 2025; Srivastava and Yao, 2025; Dutt et al., 2025), carefully designing how hierarchical devices in AMRs (e.g. varible names, parentheses, indents, and newlines) could be represented in the embedding space of encoder-only models may be worth further investigations.
@@
Appendix A Details on fine-tuning experimental setup
In our experiments, we adopt the setup and hyperparameters of previous work in MNLI fine-tuning and HANS evaluation (McCoy et al., 2020; Min et al., 2020). We take the bert-base-uncased model and fine-tune for 3 epochs with a learning rate of 2e-5. While we opt to follow prior work, we note that longer fine-tuning beyond 10 epochs at the same learning rate significantly improves HANS performance in all settings. Each label prediction is made from the [CLS] token’s final layer embedding. While the setup is equivalent to those from prior work, we implement the setup in a more modern, current stack. Due to updates in the hardware and software since prior work, slight changes in the resulting model weights are possible. To control for such effect, we perform a sanity check via baseline in-distribution test set evaluation.
Appendix B Prompts in Experiment 2
We use the following prompt template:
Premise: [X].
Hypothesis: [Y].
Where [X] and [Y] are replaced with the premise and hypothesis in question. The prompt applies to both zero and few-shot setting.
Appendix C Fine-tuning Error Analyses
C.1 Intensified subsequence overlap heuristic with AMR
@@
It is less likely that the model adopt the subsequence and constituency heuristic, as no text subsequences or constituencies are provided in the training dataset–only AMR parses are provided as input. However, it is possible that a new heuristic had formed. Consider the following two versions of premise-hypothesis pairs:
@@
C.2 Cross-setting evaluation analysis
@@
First, we observe that cross-evaluation models still perform above chance in MNLI evaluation (0.33), as seen in Table 3, which indicates both text and AMR knowledge can be leveraged despite noise from unseen form. The +AMR models’ single-mode (text only or AMR only) MNLI accuracies, together with the lower performance of AMR-only models compared to text-only models indicate that AMR information is more difficult to acquire and use than text input.
Second, we observe that bias towards entailment or non-entailment in MNLI and HANS is strongly correlated, given train-evaluation mismatch ($m=0.6$, $R^{2}=0.87$). Cross-setting evaluation results support the case of a newly developed heuristic for AMR only models, as single-mode models overwhelmingly predict entailment in HANS examples even when evaluated on +AMR settings, both at above 96%, as seen in Table 4.
@@
Appendix D LLM prediction statistics
The results are reported in Table 5.
@@ Appendix D LLM prediction statistics
Appendix E LLM setting error analysis: Example
@@
Appendix F 5-Shot Prompting Result
We report the results of our 5-shot prompting experiments in Table 6.

### v2

@@ Abstract
Natural Language Inference (NLI) relies heavily on adequately parsing the semantic content of the premise and hypothesis. In this work, we investigate whether adding semantic information in the form of an Abstract Meaning Representation (AMR) helps pretrained language models better generalize in NLI. Our experiments [^1] We publicly release our code at https://github.com/Aatlantise/advarsarial-nli-amr. integrating AMR into NLI in both fine-tuning and prompting settings show that the presence of AMR in fine-tuning hinders model generalization while prompting with AMR leads to slight gains in GPT-4o. However, an ablation study reveals that the improvement comes from amplifying surface-level differences rather than aiding semantic reasoning. This amplification can mislead models to predict non-entailment even when the core meaning is preserved.
@@
Junghyun Min   Xiulin Yang   Shira Wein Georgetown University  Amherst College {jm3743, xy236}@georgetown.eduswein@amherst.edu
@@ 1 Introduction
Natural language inference (NLI; Dagan et al., 2010) is a popular task in NLP where the solver is given a premise and a hypothesis, and asked to determine whether the hypothesis is true if the premise is true. The label space consists of three labels: entailment if the hypothesis is true, contradiction if the hypothesis is false, and neutral if the truth value of the hypothesis cannot be determined; this can also be condensed in two labels: entailment and non-entailment. As shown in Figure˜1 “Athletes introduced the secretaries” should be entailed by “Serious athletes introduced the secretaries.” Therefore, the label should be entailment because the truth of the premise indicates truth of (or entails) the hypothesis.
As a meaning-focused task, NLI aligns well with the motivation behind AMRs, i.e., to abstract sentence meaning beyond surface form, given NLI models’ tendencies to adopt shallow heuristics rather than understanding the relationship between the premise and the hypothesis, leading to poor generalization to novel data (Gururangan et al., 2018; Poliak et al., 2018; McCoy et al., 2019; Serrano et al., 2023). In this paper, we investigate whether incorporating AMRs as additional input - either during (a) fine-tuning or (b) prompting - can encourage models to attend more to abstract meaning, thereby improving generalization and overall performance. As illustrated in Figure˜1, we add AMRs to either the training data or prompts then evaluate how the addition of AMR affects generalization performance. We find that AMRs generally hinder performance in both fine-tuning and prompting settings, with the exception of prompting on HANS. However, this improvement appears to stem from AMRs amplifying surface-level differences rather than capturing deeper semantic meaning.
@@ 2 Related Work
Several prior approaches have incorporated logical representations into NLI, for example by combining neural encoders with logical reasoning modules (Chen et al., 2021), training natural logic theorem provers (Abzianidze, 2020), extracting phrase correspondences via natural deduction proofs (Yanaka et al., 2018), or constraining large language models with natural logic inference patterns (Noble et al., 2025). While these works rely on task-specific inference rules or specialized proof systems, our use of AMRs differs in that AMRs provide a broad, task-agnostic semantic abstraction without requiring dedicated engineering. Once an AMR parser is available, AMRs can be used as direct inputs to pretrained models such as BERT (Section 3.2 and ChatGPT (Section 3.3), enabling structured input with minimal task-specific engineering.
@@ 3.1 Data & Models
In these experiments, we use two datasets: MNLI (Williams et al., 2018) and HANS (McCoy et al., 2019). MNLI is a crowdsourced dataset, with a test set that is not available to the public. We follow prior work (Wang et al., 2018; Devlin et al., 2019) in taking one of its two developmental splits as the evaluation dataset. Specifically, we take the matched developmental set to use as our evaluation dataset. The training dataset includes 297k sentence pairs, while the evaluation set contains around 10k pairs. HANS is a template-based evaluation dataset, with 30k examples. Unlike MNLI and other NLI datasets, its label space consists of only two labels–entailment and non-entailment. We follow prior work (McCoy et al., 2020; Min et al., 2020) in collapsing the model’s neutral and contradiction predictions to the single non-entailment label when calculating evaluation metrics, to accommodate the two-class label space of HANS.
@@
We manually perform a small sanity check over a subset of generated AMRs to verify that AMR parses are acceptable, but do not perform a comprehensive quality check over the entire dataset. We observe that the AMRs produced for sentences in the HANS dataset are generally acceptable, likely benefiting from the sentences’ simple structure and short length, though certainly the generated AMRs contain noise; the sentences in MNLI are longer and more complex.
@@ 3.2 Experiment 1: Can fine-tuned models benefit from AMR in NLI?
We adopt the setup and hyperparameters of previous work in MNLI fine-tuning and HANS evaluation (McCoy et al., 2020; Min et al., 2020). We take the bert-base-uncased model and fine-tune for 3 epochs with a learning rate of 2e-5. While we opt to follow prior work, we note that longer fine-tuning beyond 10 epochs at the same learning rate significantly improves HANS performance in all settings. Each label prediction is made from the [CLS] token’s final layer embedding. While the setup is equivalent to those from prior work, we implement the setup in a more modern, current stack. Due to updates in the hardware and software since prior work, slight changes in the resulting model weights are possible. To control for such an effect, we perform a sanity check via baseline in-distribution test set evaluation. Finally, we integrate AMR into the models as text, via linearization, removing all newlines and whitespace sequences longer than length two.
@@ 3.3 Experiment 2: Can prompt-based models benefit from AMR in NLI?
In this experiment, we evaluate whether incorporating AMRs improves LLMs’ performance on NLI, on both the MNLI and HANS dataset, the latter of which remains challenging even after fine-tuning. Jin et al. (2024) find that only instruction-tuned GPT models are capable of reliably processing AMRs. We therefore restrict our evaluation to GPT-4o (Hurst et al., 2024) in zero-shot and 5-shot settings.
We use the following prompt template:
Premise: [X].
Hypothesis: [Y].
Where [X] and [Y] are replaced with the premise and hypothesis in question. The prompt applies to both zero and few-shot settings. We incorporate no additional details or explanations about the task (NLI), the datasets (MNLI and HANS), or the AMRs in our prompt, to best measure the LLM’s ability to use representations of meaning for NLI, rather than perform in-context learning. However, it is possible that the model perform better with additional context on the task, dataset, or AMRs.
@@ 4.1 Experiment 1
We report the accuracies of our fine-tuning models with and without AMRs in Table˜1. We report numbers from prior work (McCoy et al., 2020; Min et al., 2020) in addition to our experiments to serve as comparison baselines and to ensure our setup is correct. Our reported numbers are an average across 10 runs with varying seed.
As shown in Table˜1, AMR augmentation does not yield improvements in MNLI performance, nor HANS generalization. Perhaps analogously to previous data-driven attempts at improving generalization (Clark et al., 2019; Min et al., 2020; Yaghoobzadeh et al., 2021), additional AMR information as either text or graph does not affect MNLI performance. Analysis of their confusion matrices reveals AMR adds or subtracts little in terms of MNLI label decision boundary. On HANS performance, We discuss two main findings.
@@
AMR-only models predict the entailment label for 98.3% of HANS examples, compared to the baseline models at 94.7%. We attribute this to an intensified version of the baseline models’ heuristic correlating overlap between the hypothesis and premise to the entailment label, dubbed the lexical overlap heuristic (McCoy et al., 2019). We note this is concurrent with a still competitive MNLI performance, at 84%. We discuss this phenomenon in more detail in Sections˜A.1 and A.2.
@@
On the other hand, combining AMR information with text strongly affects HANS label decision boundaries in the opposite direction, overriding various shallow heuristics that favor the entailment label observed in McCoy et al. (2020) and in our baseline and AMR-only experiments. Our +AMR as text models 86.6% of HANS examples, and +AMR as graph models 86.9%, even predicting non-entailment on highly overlapping examples. We attempt to disentangle the effects of AMRs and text in a combined representation in Section˜A.2, where we find that while AMR can be used to perform NLI, it is less effective than text input and combining the two introduces new artifacts that are more difficult to interpret.
@@ 4.2 Experiment 2
The results for prompting with GPT-4o are shown in Table˜2. We report only the zero-shot results in the main text, as they yield similar overall performance and prediction patterns. Results for the five-shot setting are provided in Appendix˜D. Two main observations emerge.
@@
As shown in the table, model performance is consistently lowest when prompted with AMRs alone, while including the original sentence improves results. We find this is because AMRs lead models to make more negative predictions (see Appendix˜B). To test this statistically, we fit a logistic regression model predicting non-entailment using SMATCH++ (Opitz, 2023) between hypothesis and premise AMRs and data source (gold vs. predicted). A significant negative interaction ($\beta$ = -0.042, $p$ < 2e-16) shows that SMATCH similarity influences model predictions more than gold labels.
@@
Further analysis reveals that AMR’s sensitivity to surface-level lexical and syntactic variation leads to low structural overlap between semantically equivalent expressions, [^3] See Appendix C for an example. misleading the model toward non-entailment. This also explains why, on the HANS test set, prompts that include both the sentence and its AMR lead to the highest rate of negative predictions: the AMR representation amplifies subtle differences between two otherwise similar strings, making semantic mismatches more salient and pushing the model toward rejecting entailment. Such nuanced contrasts are what HANS is designed to probe in language models, prompting GPT-4o to overpredict non-entailment.
@@
As shown in Figure˜2, when GPT-4o is prompted with both sentence and AMR inputs, accuracy slightly increases for inputs longer than 80 words. However, this performance remains lower than that of sentence-only prompts across most length bins. We find no evidence that AMR-only prompts enhance robustness to longer sequences.
@@ 4.3 Summary
Our LLM experiments showcase similar preference towards the non-entailment label. This suggests that AMRs effectively highlight subtle distinctions between minimal pairs, explaining improved HANS performance. However, for simpler examples, this heightened contrast can cause the model to overpredict No, even for entailing sentence pairs.
@@ 5 Conclusion
Future work could explore more diverse linearization formats, prompt designs, and integration strategies that align AMR structure more directly with model attention or reasoning processes.
Encoder-based models have been shown to be sensitive to minor perturbations in input (Sinha et al., 2021; Jin et al., 2020), and prior work integrating AMR graphs into neural models have used a variety of formats (Wein and Opitz, 2024). Thus, in addition to Python-like and natural language-like representation of AMR’s structure (Srivastava et al., 2025; Srivastava and Yao, 2025; Dutt et al., 2025), carefully designing how hierarchical devices in AMRs (e.g. variable names, parentheses, indents, and newlines) could be represented in the embedding space of encoder-only models may be worth further investigations.
@@
Responsible Research Statement
We use ChatGPT-4o (Hurst et al., 2024) as a coding assistant during the implementation of our experiments, in addition to as a natural language processor.
Acknowledgments
This work was developed on a course project for Advanced Semantic Representation at Georgetown University. We thank Nathan Schneider and our classmates from the course for their helpful comments. A portion of experiments was conducted on the gcp-gu-hpc-cli cluster within the High Performance Computing at Georgetown University infrastructure.
Appendix A Fine-tuning Error Analyses
A.1 Intensified subsequence overlap heuristic with AMR
@@
It is less likely that the model adopts the subsequence and constituency heuristic, as no text subsequences or constituencies are provided in the training dataset–only AMR parses are provided as input. However, it is possible that a new heuristic had formed. Consider the following two versions of premise-hypothesis pairs:
@@
A.2 Cross-setting evaluation analysis
@@
First, we observe that cross-evaluation models still perform above chance in MNLI evaluation (0.33), as seen in Table˜3, which indicates both text and AMR knowledge can be leveraged despite noise from unseen form. The +AMR models’ single-mode (text only or AMR only) MNLI accuracies, together with the lower performance of AMR-only models compared to text-only models indicate that AMR information is more difficult to acquire and use than text input.
Second, we observe that bias towards entailment or non-entailment in MNLI and HANS is strongly correlated, given train-evaluation mismatch ($m=0.6$, $R^{2}=0.87$). Cross-setting evaluation results support the case of a newly developed heuristic for AMR only models, as single-mode models overwhelmingly predict entailment in HANS examples even when evaluated on +AMR settings, both at above 96%, as seen in Table˜4.
@@
Appendix B LLM prediction statistics
The results are reported in Table˜5.
@@ Appendix B LLM prediction statistics
Appendix C LLM setting error analysis: Example
@@
Appendix D 5-Shot Prompting Result
We report the results of our 5-shot prompting experiments in Table˜6.